1. A water pump operating under steady flow conditions delivers 33 kgm/s of

water from initial pressure of 82 kPaa to a final pressure of 275 kPaa. The

diameter of the inlet pipe is 15 cm and the diameter of the discharge pipe is

10 cm. What is the steady flow work if Q=0 and density of water is 1000

kgm/m3?


1
Expert's answer
2022-01-28T13:59:45-0500

"H=\frac{p_2-p_1}{\rho g}+\frac{v_2^2-v_1^2}{2g},"

"\dot m=\rho A_1v_1=\rho A_2v_2," "\implies"

"v_1=2.074~\frac ms,~v_2=4.667~\frac ms,"

"H=20.6~m."